Hayhayz (hayhayz):

Will fan and medal! How many moles of ethane (C2H6) would be needed to react with 62.3 grams of oxygen gas? Unbalanced equation: C2H6 + O2 → CO2 + H2O Show, or explain, all of your work along with the final answer.

OpenStudy (anonymous):

2 C2H6 +7 O2 = 4 CO2 + 6 H2O moles O2 = 62.3 g / 32 g/mol=1.95 the ratio between C2H6 and O2 is 2 : 7 moles C2H6 required = 1.95 x 2/7=0.557

Hayhayz (hayhayz):

@Austin316 thanks!! do you mind helping with a couple more?

OpenStudy (anonymous):

yeahh i can try.

Hayhayz (hayhayz):

What is the limiting reactant when 4.5 moles of aluminum react with 6.7 moles of oxygen gas? Unbalanced equation: Al+O2--> Al2O3 Show or explain all of your work along with the final answer.

OpenStudy (anonymous):

1) first balance the reaction: 4 Al + 3 O2 -----> 2 Al2O3 4 moles of aluminum reacts with 3 moles oxygen to form 2 moles aluminum oxide. 4.5 moles of aluminum would require (x 3/4)= 3.38 moles O2. Since there are 6.7 moles oxygen, aluminum is limiting reactant

Hayhayz (hayhayz):

4.5 moles of aluminum would require (x 3/4)= 3.38 moles O2. can you explain this part I'm a little confused haha